# Uncertainty-Guided Appearance-Motion Association Network for Out-of-Distribution Action Detection

### v1

@@ I Introduction
Deep neural networks (DNNs) [1, 2, 3, 4, 5, 6, 7, 8] have achieved impressive success on the multimedia sample detection task under a closed-set assumption [9, 10, 11, 12, 13, 14], where all the classes experienced during the test have been seen during training [15, 16, 17, 18, 19, 20]. However, vanilla DNN-based methods compulsorily classify each sample into some of the known classes [21, 22, 23, 24, 25, 26, 27, 28, 29]. It will lead to irrecoverable losses when classifying some outliers into wrong classes in some high-risk scenarios, such as autonomous driving [30, 31, 32, 33, 34, 35, 36]. Thus, the out-of-distribution (OOD) detection task [37, 38, 39, 40] is proposed to accurately detect these outliers from OOD classes and correctly classify the samples from in-distribution (ID) classes during testing. Although most OOD detection methods [41, 42, 43, 44, 45, 46] have achieved remarkable performance, they only focus on static images shown in Fig. 1, which limits their multimedia applications, where multimedia data is dynamically varying and not static. For example, in the autonomous driving scenario, we have to respond when an unknown object appears or moves on the road. These autonomous driving videos are dynamic, while previous OOD detection works focus on image datasets under the static assumption. It is unrealistic to first divide each video into multiple images and then detect unknown objects/actions [47, 48, 49, 50, 51]. This dynamic OOD detection problem based on videos is largely unexplored in the studies.
@@
II Related Works
@@
Temporal action detection. Temporal action detection (TAD) focuses on localizing action instances and classifying their categories in untrimmed videos. Although these TAD methods have achieved decent success, they regard all the actions in this video as belonging to some of the pre-defined classes. Besides, they only rely on the motion feature, which captures the redundant background information and fails to perceive the fine-grained differences among video frames with high similarity. For example, two real-world actions “HighJump” and “LongJump” belong to different action classes. As shown in Fig. 2(a), the main difference between “HighJump” and “LongJump” is the appearance information: “high jump bar” in “HighJump” and “sandpit” in “LongJump”. Since they only model the motion features to capture the same motion “jump”, they cannot correctly distinguish the local details of different objects (“high jump bar” and “sandpit”) in these frames, which will lead to wrong action detection results. Although some TAD methods utilize the object detection technologies for action detection, these methods focus on single-object detection setting so that they only detect one object “human” and ignore other relevant objects. Thus, these methods cannot achieve satisfactory performance, where supplementary file illustrates two qualitative results. Different from them, we creatively explore both appearance and motion features among different objects (e.g., “high jump bar” and “sandpit”) for action detection. By integrating appearance and motion features, we can correctly distinguish “HighJump” and “LongJump” via inter-object interaction.
@@
Out-of-distribution detection. As a challenging multimedia task, out-of-distribution (OOD) detection targets to detect test samples from distributions that do not overlap with the training distribution. Different from previous OOD detection works that cannot handle video datasets, our proposed method can detect OOD actions in the challenging video dataset. Therefore, our method can be applied to more challenging scenarios than these image-based OOD detection methods.
@@ III-C Uncertainty-guided OOD Action Detection
Action-background segmentation. Real-world multimedia videos naturally mix background and actions (ID and OOD actions), which makes it insufficient to distinguish them only through classification and uncertainty. Therefore, we design an action detection module to indicate the likelihood of an frame belonging to an action. For convenience, we treat the action-background segmentation as a binary classification task.
To distinguish class-agnostic actions from backgrounds, we follow [56] to introduce the frame-aware actionness score: $\hat{a}_{i}\in[0,1]$. For any two frames $x_{t_{1}}$ and $x_{t_{2}}$ ($t_{1}\neq t_{2}$ and $x_{t_{1}},x_{t_{2}}\in\{1,\cdots,T\}$), they have three relationships: 1) both $x_{t_{1}}$ and $x_{t_{2}}$ belong to the action; 2) both $x_{t_{1}}$ and $x_{t_{2}}$ are from the background; 3) one belongs to the action, while the other is from the background. Therefore, we introduce the cosine similarity $\cos(x_{t_{1}},x_{t_{2}})$ and three thresholds ($\tau_{bb},\tau_{aa},\tau_{dif}$) to evaluate their relationship. Based on the annotated background frames and potential action frames, we can calculate the affinity loss $\mathcal{L}_{ABS}$ from three terms, i.e., between two background frames $\mathcal{L}^{\rm bg}$, between two action frames $\mathcal{L}^{\rm act}$ and between the action-background pair $\mathcal{L}^{\rm dif}$.
Particularly, we use the online hard example mining strategy [57] to constrain the training frame pair. We align two background frames into the same class by the following loss:
@@
where $a_{\tau}$ is an actionness threshold; $\tau_{bb}$ is the similarity threshold between two frames from the background class. Also, we utilize a similar loss to align two action frames:
@@
where $\tau_{aa}$ is the similarity threshold between two frames from the action class. When $x_{t_{1}}$ and $x_{t_{2}}$ belong to action and background respectively, we can obtain the following loss:
@@
where $\tau_{dif}$ is the threshold to constrain the similarity between actions and backgrounds. The affinity loss jointly considers the above three terms and can be calculated as:
@@
Action classification. Different from previous OOD detection methods, we are required to detect OOD actions in an untrimmed video. In real-world multimedia scenario, the predicted likelihood for each action class follows a binomial distribution whose conjugate prior is a Beta distribution. We introduce evidential neural networks [58] based on Beta distribution to jointly formulate the multi-class classification and uncertainty modeling. We assume a Beta distribution $Beta(p_{i}|\alpha_{i},\beta_{i})$ over the action categorical probability $p_{i}\in[0,1]$, where $\alpha_{i}$ and $\beta_{i}$ are parameters indicating the positive and negative evidence of the $i$-th action, respectively. The evidence evaluates actions closest to the predicted ones in the feature space and is used to support the decision-making. Predicted labels should be the same as positive evidence, but different from negative evidence.
Based on $\alpha_{i}$ and $\beta_{i}$, we can design a subjective opinion $w_{i}=(b_{i},d_{i},u_{i},a_{i})$ for action $i\in\{1,2,\cdots,K\}$, where belief $b_{i}\in[0,1]$, disbelief $d_{i}\in[0,1]$, uncertainty score $u_{i}\in[0,1]$, and $b_{i}+d_{i}+u_{i}=1$. For action $i$, the opinion $w_{i}$ is obtained by $\alpha_{i}$ and $\beta_{i}$: $b_{i}=\frac{\alpha_{i}-2}{\alpha_{i}+\beta_{i}},\quad d_{i}=\frac{\beta_{i}-2}{\alpha_{i}+\beta_{i}},\quad u_{i}=\frac{2}{\alpha_{i}+\beta_{i}}$. The expected belief probability $p_{i}=b_{i}+a_{i}\cdot u_{i}$, where $a_{i}$ refers to a base rate representing prior knowledge without commitment, such as neither agree nor disagree.
For an object $j$, its positive and negative evidence are estimated as $\boldsymbol{\alpha}_{j}=s(h(\mathbf{x}_{j};\boldsymbol{\theta}))+1$ and $\boldsymbol{\beta}_{j}=s(h(\mathbf{x}_{j};\boldsymbol{\theta}))+1$, where $\boldsymbol{\alpha}_{j}=[\alpha_{1j},\cdots,\alpha_{Kj}]^{\top}$ and $\boldsymbol{\beta}_{j}=[\beta_{1j},\cdots,\beta_{Kj}]^{\top}$. $\mathbf{x}_{j}$ denotes the input video, $h(\mathbf{x}_{j};\boldsymbol{\theta})$ represents the evidence vector predicted by the network for the classification and, $\boldsymbol{\theta}$ represents parameters for modeling. Here $s(\cdot)$ is the evidence function (e.g., ReLU) to keep $\boldsymbol{\alpha}_{j},\boldsymbol{\beta}_{j}\geq\mathbf{1}$.
To learn the above opinions, we define the Beta loss function by computing its Bayes risk for the action predictor. For the binary cross-entropy loss for each action $i$ over a batch of actors, we introduce the following loss:
@@
where $K$ is the number of actions, and $BC(\cdot)$ is the binary cross-entropy loss, and $\psi(\cdot)$ is the digamma function. The log expectation of Beta distribution derives the last equality. $\mathbf{y}_{j}=[y_{1j},\cdots,y_{Kj}]\in\{0,1\}^{K}$ is the $K$-dimensional ground-truth action label for $\mathbf{x}_{j}$.

### v2

@@ I Introduction
Deep neural networks (DNNs) [1, 2, 3, 4, 5, 6, 7, 8] have achieved impressive success on the multimedia sample detection task under a closed-set assumption [9, 10, 11, 12, 13, 14, 15, 16, 17, 18], where all the classes experienced during the test have been seen during training [19, 20, 21, 22, 23, 24]. However, vanilla DNN-based methods compulsorily classify each sample into some of the known classes [25, 26, 27, 28, 29, 30, 31, 32, 33]. It will lead to irrecoverable losses when classifying some outliers into wrong classes in some high-risk scenarios, such as autonomous driving [34, 35, 36, 37, 38, 39, 40]. Thus, the out-of-distribution (OOD) detection task [41, 42, 43, 44] is proposed to accurately detect these outliers from OOD classes and correctly classify the samples from in-distribution (ID) classes during testing. Although most OOD detection methods [45, 46, 47, 48, 49, 50] have achieved remarkable performance, they only focus on static images shown in Fig. 1, which limits their multimedia applications, where multimedia data is dynamically varying and not static. For example, in the autonomous driving scenario, we have to respond when an unknown object appears or moves on the road. These autonomous driving videos are dynamic, while previous OOD detection works focus on image datasets under the static assumption. It is unrealistic to first divide each video into multiple images and then detect unknown objects/actions [51, 52, 53, 54, 55]. This dynamic OOD detection problem based on videos is largely unexplored in the studies.
@@
II Related Work
Temporal action detection. Temporal action detection (TAD) [59, 60, 61] focuses on localizing action instances and classifying their categories in untrimmed videos [62, 63, 15]. Unlike traditional action detection [64, 65, 66, 67] that focuses on identifying actions in videos without specifying their temporal boundaries, TAD involves precisely localizing the start and end times of the target actions within an untrimmed video. Based on existing action detection models, OpenMax [68] generates the action proposals, and employs OpenMax in testing to append the softmax scores with OOD actions. DEAR [69] replaces the traditional cross-entropy loss for uncertainty quantification to detect OOD actions. By predicting the locations, classifications with uncertainties, and actionness, OpenTAL recognizes and localizes all the actions simultaneously, and rejects OOD actions. TFE-DCN [70] designs a temporal feature enhancement dilated convolution network for weakly-supervised action detection. Recently, some methods have been proposed based on the detection framework [71, 72, 73]. For example, [71] aims to directly predict the boundary of target actions by designing an end-to-end framework. Although these TAD methods have achieved decent success, they regard all the actions in this video as belonging to some of the pre-defined classes.
@@ II Related Work
Unfortunately, they only rely on the motion feature, which captures the redundant background information and fails to perceive the fine-grained differences among video frames with high similarity. For example, two real-world actions “HighJump” and “LongJump” belong to different action classes. As shown in Fig. 2, the main difference between “HighJump” and “LongJump” is the appearance information: “high jump bar” in “HighJump” and “sandpit” in “LongJump”. Since they only model the motion features to capture the same motion “jump”, they cannot correctly distinguish the local details of different objects (“high jump bar” and “sandpit”) in these frames, which will lead to wrong action detection results. Although some TAD methods [74, 63] utilize the object detection technologies for action detection, these methods focus on single-object detection setting so that they only detect one object “human” and ignore other relevant objects. Thus, these methods cannot achieve satisfactory performance, where Fig. LABEL:fig:vis_demo illustrates two qualitative results. Different from them, we creatively explore both appearance and motion features among different objects (e.g., “high jump bar” and “sandpit”) for action detection. By integrating appearance and motion features, we can correctly distinguish “HighJump” and “LongJump”.
Out-of-distribution detection. As a challenging computer vision task, out-of-distribution (OOD) detection targets to detect test samples from distributions that do not overlap with the training distribution. Previous OOD detection methods [45, 46, 47, 48, 49, 75, 76, 77, 42, 50] can be divided into four types: classification-based methods [78, 45, 79, 80], density-based methods [81, 82], distance-based methods [83, 48] and reconstruction-based methods [84, 85]. 1) Early OOD detection works refer to a classification framework, which utilizes the maximum softmax probability to determine the ID/OOD samples. 2) To more explicitly model ID, density-based OOD detection methods are proposed to leverage the probabilistic models for OOD detection. These methods are under an operating assumption that OOD samples have low likelihoods whereas ID samples have high likelihoods under the estimated density model. 3) The distance-based OOD detection methods are based on an intuitive idea that OOD samples should be relatively far away from the centroids of ID samples. under the Linear Discriminant Analysis assumption [86], and then utilizes the minimum Mahalanobis distance to all class centroids for OOD detection. 4) The reconstruction-based methods often leverage the encoder-decoder framework, which is trained on only ID samples and generates different outcomes for OOD detection.
Different from previous OOD detection works that cannot handle video datasets, our proposed method can detect OOD actions in the challenging video dataset. Therefore, our method can be applied to more challenging scenarios than these image-based OOD detection methods.
@@ III-C Uncertainty-guided OOD Action Detection
Action-background segmentation. Real-world videos naturally mix background and actions (ID and OOD actions), making it insufficient to recognize them only based on classification and uncertainty. Therefore, we design an action detection module to indicate the likelihood of an frame belonging to an action. For convenience, we treat the action-background segmentation as a binary classification task.
To recognize class-agnostic actions from background frames, we follow [88] to introduce the frame-aware actionness score: $\hat{a}_{i}\in[0,1]$. For any two frames $x_{t_{1}}$ and $x_{t_{2}}$ ($t_{1}\neq t_{2}$ and $x_{t_{1}},x_{t_{2}}\in\{1,\cdots,T\}$), they have three relationships: 1) both $x_{t_{1}}$ and $x_{t_{2}}$ belong to the action; 2) both $x_{t_{1}}$ and $x_{t_{2}}$ are from the background; 3) one belongs to the action, while the other is from the background. Therefore, we introduce the cosine similarity $\cos(x_{t_{1}},x_{t_{2}})$ and three thresholds ($\tau_{bb},\tau_{aa},\tau_{dif}$) to evaluate their relationship. To supervise the action-background segmentation module by the annotated background frames and potential action frames, we introduce an affinity loss $\mathcal{L}_{ABS}$. Especially, $\mathcal{L}_{ABS}$ contains three significant parts ($\mathcal{L}^{\rm bg}$ based on two background frames, $\mathcal{L}^{\rm act}$ based on two action frames, and $\mathcal{L}^{\rm dif}$ based on the action-background pair).
For $\mathcal{L}^{\rm bg}$, we use the online hard example mining strategy [89] to constrain two training background frames. We align two background frames into the same class by the following loss:
@@
where $a_{\tau}$ is an actionness threshold; $\tau_{bb}$ is the similarity threshold between two background frames from the background class. Also, we utilize a similar loss to align two action frames:
@@
where $\tau_{aa}$ is the similarity threshold between two action frames from the action class. When $x_{t_{1}}$ and $x_{t_{2}}$ belong to action and background respectively, we can obtain the following loss:
@@
where $\tau_{dif}$ is the threshold to constrain the similarity between an action frame and a background frame. Based on the above three parts ($\mathcal{L}^{\rm bg}$, $\mathcal{L}^{\rm act}$ and $\mathcal{L}^{\rm dif}$), we can obtain the affinity loss $\mathcal{L}_{ABS}$ as follows:
@@
Action classification. Different from previous OOD detection methods, we are required to detect OOD actions in an untrimmed video. For an action class in real-world multimedia scenario, its predicted likelihood always follows a binomial distribution, and its conjugate prior is treated as a Beta distribution. We introduce evidential neural networks [90] based on Beta distribution to jointly formulate the multi-class classification and uncertainty modeling. For the $i$-th action, we assume a Beta distribution $Beta(p_{i}|\alpha_{i},\beta_{i})$ over the action categorical probability $p_{i}\in[0,1]$, where $\alpha_{i}$ and $\beta_{i}$ denote parameters to indicate the positive and negative evidence, respectively. In the feature space, the evidence evaluates actions closest to the predicted actions. Predicted labels should be the same as positive evidence, but different from negative evidence.
For action $i\in\{1,2,\cdots,K\}$, we can design a subjective opinion $w_{i}=(b_{i},d_{i},u_{i},a_{i})$ based on $\alpha_{i}$ and $\beta_{i}$, where belief $b_{i}\in[0,1]$, disbelief $d_{i}\in[0,1]$, uncertainty score $u_{i}\in[0,1]$, and $b_{i}+d_{i}+u_{i}=1$. For action $i$, the opinion $w_{i}$ is obtained by $\alpha_{i}$ and $\beta_{i}$: $b_{i}=\frac{\alpha_{i}-2}{\alpha_{i}+\beta_{i}},\quad d_{i}=\frac{\beta_{i}-2}{\alpha_{i}+\beta_{i}},\quad u_{i}=\frac{2}{\alpha_{i}+\beta_{i}}$. The expected belief probability $p_{i}=b_{i}+a_{i}\cdot u_{i}$, where $a_{i}$ is a base rate that denotes prior knowledge without commitment (neither agree nor disagree).
For object $j$, its positive evidence is $\boldsymbol{\alpha}_{j}=s(h(\mathbf{x}_{j};\boldsymbol{\theta}))+1$ and its negative evidence is $\boldsymbol{\beta}_{j}=s(h(\mathbf{x}_{j};\boldsymbol{\theta}))+1$, where $\boldsymbol{\alpha}_{j}=[\alpha_{1j},\alpha_{2j},\cdots,\alpha_{Kj}]^{\top}$ and $\boldsymbol{\beta}_{j}=[\beta_{1j},\beta_{2j},\cdots,\beta_{Kj}]^{\top}$. $\mathbf{x}_{j}$ means the video, $h(\mathbf{x}_{j};\boldsymbol{\theta})$ means the evidence vector and $\boldsymbol{\theta}$ mean parameters, $s(\cdot)$ means the evidence function (e.g., ReLU) to guarantee $\boldsymbol{\alpha}_{j},\boldsymbol{\beta}_{j}\geq\mathbf{1}$.
For learning the above opinions, we introduce the Beta loss function by computing its Bayes risk for our action classification module. For the binary cross-entropy loss for each action $i$ over a batch of actors, we introduce the following loss:
@@
where $K$ denotes the action number, $BC(\cdot)$ denotes the binary cross-entropy loss, and $\psi(\cdot)$ denotes the digamma function. The log expectation of Beta distribution derives the last equality. For video $\mathbf{x}_{j}$, $\mathbf{y}_{j}=[y_{1j},y_{2j},\cdots,y_{Kj}]\in\{0,1\}^{K}$ denotes its $K$-dimensional ground-truth action label.